# Bench2Drive: Towards Multi-Ability Benchmarking of Closed-Loop End-To-End Autonomous Driving

### v2

@@ Abstract
To fulfill the paramount need of comprehensive, realistic, and fair testing environments for Full Self-Driving (FSD), we present Bench2Drive, the first benchmark for evaluating E2E-AD systems’ multiple abilities in a closed-loop manner. Bench2Drive’s official training data consists of 2 million fully annotated frames, collected from 10000 short clips uniformly distributed under 44 interactive scenarios (cut-in, overtaking, detour, etc), 23 weathers (sunny, foggy, rainy, etc), and 12 towns (urban, village, university, etc) in CARLA v2. Its evaluation protocol requires E2E-AD models to pass 44 interactive scenarios under different locations and weathers which sums up to 220 routes and thus provides a comprehensive and disentangled assessment about their driving capability under different situations. We implement state-of-the-art E2E-AD models and evaluate them in Bench2Drive, providing insights regarding current status and future directions.
@@ 1 Introduction
In recent years, the field of autonomous driving has witnessed tremendous growth, fueled by the rapid advancement and scaling of foundation models [2, 37, 51]. These developments have ushered in a new era of end-to-end autonomous driving (E2E-AD) systems [18, 48, 8], which promise a scalable, data-driven approach to vehicle automation, opposed to traditional module-based perception [31, 33, 35], prediction [22, 23, 25, 27], planning [21, 11] pipeline. Such systems are designed to be capable of learning from vast amounts of data, potentially transforming the landscape of vehicle intelligence.
@@
CARLA [13] is one of the most widely used simulator for closed-loop E2E-AD evaluation. Within its framework, benchmarks such as Town05Long and Longest6 have been established, featuring multiple routes that require AD systems to complete safely within specific time constraints. However, these benchmarks only assess basic skills such as lane following, making turns, collision avoidance, and traffic lights obeying [5, 26], failing to examine AD systems’ driving ability under complicated and interactive traffic. The latest CARLA Leaderboard v2 introduces 39 challenging scenarios designed to evaluate the robustness of AD systems in more intricate situations. Nevertheless, the official routes for evaluation, ranging from 7 to 10 kilometers and filled with scenarios, present a formidable challenge, often too difficult to complete flawlessly, as shwon in Fig. 2 (a). Consequently, with the driving score metric employing an exponential decay function, it becomes challenging to effectively compare different AD systems, as they tend to score very low. For instance, in the current Leaderboard v2111https://eval.ai/web/challenges/challenge-page/2098/leaderboard/4942, participating methods score less than 10 points out of 100. Besides, existing methods usually collect data by themselves which makes algorithm-level fair comparison infeasible.
@@
To address the aforementioned challenges in evaluating autonomous driving (AD) systems, it is essential to develop a new benchmark that fairly assesses their capabilities in a granular manner. To this end, we introduce Bench2Drive, a new benchmark designed to evaluate E2E-AD systems in a comprehensive, realistic, and fair closed-loop environment. Bench2Drive has an official training dataset collected by state-of-the-art epxert model Think2Drive [32], comprising 2 million fully annotated frames, sourced from 10000 clips. It span a diverse array of 44 interactive scenarios such as cut-ins, overtakings, and detours under different weather conditions and towns, ranging from sunny days in bustling city centers to foggy conditions in quaint villages. The evaluation protocol includes 220 short routes, each only around 150 meters in length and containing a single specific scenario. In this way, the assessment of individual skills is isolated and thus allows for a detailed comparison of the AD systems’ proficiency across 44 distinct skill sets. Moreover, the brevity of each route mitigates the impact of the exponential decay function on the driving score, facilitating a more accurate and meaningful comparison of performance across different systems. Such a structured and focused benchmark would provide clearer insights into the strengths and weaknesses of each AD system, enabling targeted improvements and more refined technology development.
@@
  • Diverse Large-Scale Official Training Data: Bench2Drive consists of a standardized training set of 2 million fully annotated frames from 10000 clips under diverse scenarios, weathers, and towns, ensuring that all AD systems are trained under abundant yet similar conditions, which is crucial for fair algorithm-level comparisons.
These features make Bench2Drive a pioneering benchmark in the field of autonomous driving, providing an essential tool for researchers to refine and evaluate their E2E-AD systems in a realistic, comprehensive, and fair manner. We implement several classic baselines including TCP [48], ThinkTwice [26], DriveAdapter [24], UniAD [18], VAD [28], and AD-MLP [52] and evaluate them in the Bench2Drive. We confirm the fact that open-loop metrics like L2 error could not reflect the actual driving performance. For the classic clsoed-loop metric - Drive Score, we find that it lack details and its heavy punishment encourages over-conservative driving strategies while Bench2Drive offers a comprehensive understanding about capabilities of different methods.
@@ 2.1 Planning Benchmarks
Benchmarking in the field of autonomous driving has evolved from specialized datasets, such as KITTI [14] for perception and NGSIM/highD [30], BARK [1] for behavior prediction, to integrated forms like nuScenes [3], Argoverse [47], and Waymo [44], which facilitate the evaluation of various synergic system components. Recently, the assessment of planning capabilities for learning-based methods has become an area of interest [15, 29]. In Table 1, we present a comparison of planning benchmarks. nuScenes [3], while offering open-loop metrics, has been critiqued for its inability to adequately evaluate planning proficiency due to the lack of closed-loop simulation [52, 34, 11]. Furthermore, it suffers from an imbalanced validation set, with a significant portion (75%) of scenarios only requiring straightforward driving, thus inadequately challenging the decision-making capabilities of AD systems in complex environments [34]. nuPlan [29] and Waymax [15] offer closed-loop evaluations but are limited to bounding box level assessments, excluding sensor simulation and, consequently, are not suitable for E2E-AD methods. Longest6 [8], a modified version of CARLA Leaderboard V1, only assesses basic skills such as lane following, making turns, collision avoidance, and traffic lights. CARLA Leaderboard V2 [13] lacks expert demonstration data. As widely discussed in the community [40, 19], the lack of an official training set makes the comparisons of different methods in the system-level instead of the algorithm-level.Bench2Drive deal with these shortcomings by offering a large-scale, annotation-rich official training dataset alongside a multi-ability evaluation set. This enables a more granular and informative assessment of an AD system’s driving capabilities, overcoming the limitations of existing benchmarks that rely on average scoring across all routes as their primary performance metric.
Table 1: Comparison with related planning benchmarks.
@@ 3.1 Expert Dataset
Figure 4: Distribution of scenario ’ConstructionObstacle’ in Town12
Figure 5: Scenario Distribution of Bench2Drive Dataset
Moreover, to tackle the challenge posed by the long-tail distribution of data from both perception and behavior perspectives, a significant bottleneck in autonomous driving [20] (approximately 75% of the clips in nuScenes only involve the ego vehicle driving straight), we ensure the distribution of weather conditions, landscapes, and behaviors are as uniform as possible. We add more available locations for scenarios compared to the official routes of CARLA Leaderboard V2 as shown in Fig. 4, enhancing the data diversity. Further, we design 5 more scenarios beyond Leaderboard V2 to enhance behavior diversity as detailed in Appendix D. We give the distribution of scenarios in Fig. 5 and distribution of weathers and towns in Appendix B. As illustrated, Bench2Drive dataset is rich in both perception and behavior diversity.
@@ 4.1 Baselines & Datasets
Table 3: Open-loop and Closed-loop Results of E2E-AD Methods in Bench2Drive. Avg. L2 is averaged over the predictions in 2 seconds under 2Hz. * denotes expert feature distillation.
Method Input Open-loop Metric Closed-loop Metric
Avg. L2 (meter) $\downarrow$ Driving Score $\uparrow$ Success Rate(%) $\uparrow$
AD-MLP [52] Ego State 3.64 9.14 0.00
UniAD-Tiny [18] Ego State + 6 Camera 0.80 32.00 9.54
UniAD-Base [18] Ego State + 6 Camera 0.73 37.72 9.54
VAD [28] Ego State + 6 Camera 0.91 39.42 10.00
TCP* [48] Ego State + Front Cameras 1.70 23.63 7.72
TCP-ctrl* Ego State + Front Cameras - 18.63 5.45
TCP-traj* Ego State + Front Cameras 1.70 36.78 26.82
ThinkTwice* [26] Ego State + 6 Cameras 0.95 39.88 28.14
DriveAdapter* [24] Ego State + 6 Cameras 1.01 42.91 30.71
Table 4: Multi-Ability Results of E2E-AD Methods. * denotes expert feature distillation.
@@
AD-MLP [52] 0.00 0.00 0.00 0.00 0.00 0.00
UniAD-Tiny [18] 4.11 12.50 14.54 10.00 18.54 11.94
UniAD-Base [18] 9.46 12.50 20.00 30.00 23.03 19.00
VAD [28] 12.50 17.50 14.54 30.00 25.55 20.02
TCP* [48] 8.70 10.00 7.27 10.00 7.95 8.78
TCP-ctrl* 8.57 10.00 3.63 0.00 7.95 6.03
TCP-traj* 24.29 15.00 29.09 50.00 51.67 34.01
ThinkTwice* [26] 26.44 17.50 32.12 50.00 53.65 35.94
DriveAdapter* [24] 29.23 20.00 34.71 50.00 57.21 38.23
@@ 4.2 Results
Open-loop metric could indicate model convergence but it fails for advanced comparison.. AD-MLP has a high L2 error and performs extremely bad in closed-loop evaluation while VAD has a low L2 error and a decent closed-loop performance. It shows that we could use L2 error to verify the convergence and fitting status of neural networks, i.e., when the L2 error is very high, there should be something wrong within the system. In this case, AD-MLP does not use raw sensors, which is similar to drive blindly and thus infeasible to fit the dataset. Notably, different from findings in nuScenes [3], AD-MLP fails to achieve decent L2 error in Bench2Drive, due to the better behavior diversity as shown in Fig. 6. On the other hand, UniAD-base has a lower L2 error compared to VAD yet with worse closed-loop performance, aligning with findings in [11, 34]. Open-loop evaluation ignore the issues including distribution shift [41] and causal confusion [12, 24] and thus fails to give meaningful comparsion for models with good fitting of dataset, demonstrating the importance of closed-loop evaluation.
Expert feature distillation offers important guidance. As pointed out in [6, 55], due to the high-dimensional input space of AD, i.e., multiple images and point clouds, E2E-AD methods tend to overfit. The features from expert, which already possesses strong driving knowledge, could be helpful to mitigate the issue by distillation. As a result, methods (TCP/ThinkTwice/DriveAdapter) with expert feature distillation outperforms those without (VAD/UniAD) by a large margin. However, in the real world setting, it could be difficult to obtain expert features, which worths further study.

### v3

@@ Abstract
To fulfill the paramount need of comprehensive, realistic, and fair testing environments for Full Self-Driving (FSD), we present Bench2Drive, the first benchmark for evaluating E2E-AD systems’ multiple abilities in a closed-loop manner. Bench2Drive’s official training data consists of 2 million fully annotated frames, collected from 13638 short clips uniformly distributed under 44 interactive scenarios (cut-in, overtaking, detour, etc), 23 weathers (sunny, foggy, rainy, etc), and 12 towns (urban, village, university, etc) in CARLA v2. Its evaluation protocol requires E2E-AD models to pass 44 interactive scenarios under different locations and weathers which sums up to 220 routes and thus provides a comprehensive and disentangled assessment about their driving capability under different situations. We implement state-of-the-art E2E-AD models and evaluate them in Bench2Drive, providing insights regarding current status and future directions.
@@ 1 Introduction
footnotetext: This work was in part supported by by NSFC (92370201, 62222607) and Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102.
In recent years, the field of autonomous driving has witnessed tremendous growth, fueled by the rapid advancement and scaling of foundation models [1, 2, 3]. These developments have ushered in a new era of end-to-end autonomous driving (E2E-AD) systems [4, 5, 6, 7, 8], which promise a scalable, data-driven approach to vehicle automation, opposed to traditional module-based perception [9, 10, 11, 12, 13], prediction [14, 15, 16, 17], planning [18, 19, 20] pipeline. Such systems are designed to be capable of learning from vast amounts of data, potentially transforming the landscape of vehicle intelligence.
@@
CARLA [28] is one of the most widely used simulator for closed-loop E2E-AD evaluation. Within its framework, benchmarks such as Town05Long and Longest6 have been established, featuring multiple routes that require AD systems to complete safely within specific time constraints. However, these benchmarks only assess basic skills such as lane following, making turns, collision avoidance, and traffic lights obeying [29, 30], failing to examine AD systems’ driving ability under complicated and interactive traffic. The latest CARLA Leaderboard v2 introduces 39 challenging scenarios designed to evaluate the robustness of AD systems in more intricate situations. Nevertheless, the official routes for evaluation, ranging from 7 to 10 kilometers and filled with scenarios, present a formidable challenge, often too difficult to complete flawlessly, as shown in Fig. 2 (a). Consequently, with the driving score metric employing an exponential decay function, it becomes challenging to effectively compare different AD systems, as they tend to score very low. For instance, in the current Leaderboard v2111https://eval.ai/web/challenges/challenge-page/2098/leaderboard/4942, participating methods score less than 10 points out of 100. Besides, existing methods usually collect data by themselves which makes algorithm-level fair comparison infeasible.
@@
To address the aforementioned challenges in evaluating autonomous driving (AD) systems, it is essential to develop a new benchmark that fairly assesses their capabilities in a granular manner. To this end, we introduce Bench2Drive, a new benchmark designed to evaluate E2E-AD systems in a comprehensive, realistic, and fair closed-loop environment. Bench2Drive has an official training dataset collected by state-of-the-art expert model Think2Drive [31], comprising 2 million fully annotated frames, sourced from 13638 clips. It span a diverse array of 44 interactive scenarios such as cut-ins, overtakings, and detours under different weather conditions and towns, ranging from sunny days in bustling city centers to foggy conditions in quaint villages. The evaluation protocol includes 220 short routes, each only around 150 meters in length and containing a single specific scenario. In this way, the assessment of individual skills is isolated and thus allows for a detailed comparison of the AD systems’ proficiency across 44 distinct skill sets. Moreover, the brevity of each route mitigates the impact of the exponential decay function on the driving score, facilitating a more accurate and meaningful comparison of performance across different systems. Such a structured and focused benchmark would provide clearer insights into the strengths and weaknesses of each AD system, enabling targeted improvements and more refined technology development.
@@
  • Diverse Large-Scale Official Training Data: Bench2Drive consists of a standardized training set of 2 million fully annotated frames from 13638 clips under diverse scenarios, weathers, and towns, ensuring that all AD systems are trained under abundant yet similar conditions, which is crucial for fair algorithm-level comparisons.
These features make Bench2Drive a pioneering benchmark in the field of autonomous driving, providing an essential tool for researchers to refine and evaluate their E2E-AD systems in a realistic, comprehensive, and fair manner. We implement several classic baselines including TCP [5], ThinkTwice [30], DriveAdapter [27], UniAD [4], VAD [22], and AD-MLP [23] and evaluate them in the Bench2Drive. We confirm the fact that open-loop metrics like L2 error could not reflect the actual driving performance. For the classic closed-loop metric - Drive Score, we find that it lack details and its heavy punishment encourages over-conservative driving strategies while Bench2Drive offers a comprehensive understanding about capabilities of different methods.
@@ 2.1 Planning Benchmarks
Benchmarking in the field of autonomous driving has evolved from specialized datasets, such as KITTI [32] for perception and NGSIM/highD [33], BARK [34] for behavior prediction, to integrated forms like nuScenes [21], Argoverse [35], and Waymo [36], which facilitate the evaluation of various synergic system components. Recently, the assessment of planning capabilities for learning-based methods has become an area of interest [37, 38, 39, 40, 41]. In Table 1, we present a comparison of planning benchmarks. nuScenes [21], while offering open-loop metrics, has been critiqued for its inability to adequately evaluate planning proficiency due to the lack of closed-loop simulation [23, 24, 19]. Furthermore, it suffers from an imbalanced validation set, with a significant portion (75%) of scenarios only requiring straightforward driving, thus inadequately challenging the decision-making capabilities of AD systems in complex environments [24]. nuPlan [38] and Waymax [37] offer closed-loop evaluations but are limited to bounding box level assessments, excluding sensor simulation and, consequently, are not suitable for E2E-AD methods. Longest6 [6], a modified version of CARLA Leaderboard V1, only assesses basic skills such as lane following, making turns, collision avoidance, and traffic lights. CARLA Leaderboard V2 [28] lacks expert demonstration data. As widely discussed in the community [42, 43], the lack of an official training set makes the comparisons of different methods in the system-level instead of the algorithm-level.Bench2Drive deal with these shortcomings by offering a large-scale, annotation-rich official training dataset alongside a multi-ability evaluation set. This enables a more granular and informative assessment of an AD system’s driving capabilities, overcoming the limitations of existing benchmarks that rely on average scoring across all routes as their primary performance metric.
Table 1: Comparison with related planning benchmarks Bench2Drive is the only benchmark to evaluate the E2E-AD methods under closed-loop with multi-ability analysis.
@@ 3 Bench2Drive
3.1 Data Collection Agent
The data collection agent (expert) is responsible for collecting the data so that student models could learn from the data. In the real world, this is usually done by human to drive around the city, like the curation of KITTI [32], nuScenes [21], Waymo [36], Argoverse [35]. However, it requires lots of human efforts. In simulation, there is a cheap substitute - teacher model. The teacher model would use information not available in the real world (termed privileged information), for example, ground-truth locations, states, and intentions of surrounding agents and ground-truth states of traffic lights, etc. As a result, people using CARLA either write rules [43, 61] or train a RL model [50, 31] to use the privileged information to drive in the simulation.
In this work, we use the world model based reinforcement learning teacher - Think2Drive [31] to navigate in CARLA and collect data, since it is the only expert model which is able to solve all 44 scenarios during the construction of Bench2Drive. Notably, after the release of Bench2Drive, the rule-based expert PDM-Lite [61] [^2] https://github.com/autonomousvision/carla_garage/tree/leaderboard_2 is open sourced and users could use it for customized demand.
@@ 3.2 Expert Dataset
Figure 4: Distribution of scenario ’ConstructionObstacle’ in Town12 We use different colors to represent different routes containing ’ConstructionObstacle’. Bench2Drive has more locations that be able to generate ’ConstructionObstacle’.
Moreover, to tackle the challenge posed by the long-tail distribution of data from both perception and behavior perspectives, a significant bottleneck in autonomous driving [62] (approximately 75% of the clips in nuScenes only involve the ego vehicle driving straight), we ensure the distribution of weather conditions, landscapes, and behaviors are as uniform as possible. We add more available locations for scenarios compared to the official routes of CARLA Leaderboard V2 as shown in Fig. 4, enhancing the data diversity. Further, we design 5 more scenarios beyond Leaderboard V2 to enhance behavior diversity as detailed in Appendix G. We give the distribution of scenarios and weathers and towns in Appendix B. As illustrated, Bench2Drive dataset is rich in both perception and behavior diversity.
@@ 3.3 Multi-Ability Evaluation
Further, beyond the the goal achieving ability of algorithms, we propose the following two metrics to measure the efficiency and smoothness of driving trajectories:
  • Efficiency: The CARLA team has implemented a function to check whether the self-driving car’s speed is too low. This is determined by comparing the vehicle’s speed with nearby vehicle:
    Speed Percentage=Ego Vehicle’s SpeedAverage Speed of Nearby $\text{Speed Percentage}=\frac{\text{Ego Vehicle's Speed}}{\text{Average Speed of Nearby Vehicles}}$ (2)
    This function calculates the speed percentage using the vehicle’s speed and the average speed of nearby vehicles at current frame. CARLA Leaderboard sets four checkpoints per route and checks the ego vehicle’s speed when the ego vehicle arrives a checkpoint. Specifically, if the vehicle is faster than nearby vehicles, the driving efficiency would be larger than 100%. The check results are included as a penalty in the final driving score. However, with only four checkpoints, the vehicle must cover 25% of the total route distance before reaching the next checkpoint. This leads to a high variance in the penalty values for low speeds, complicating the reflection of driving capabilities in the driving scores. To alleviate this, we increase the number of checkpoints to 20. Speed check is now performed every 5% of the total route length, and it is excluded from the driving score calculation. The final driving efficiency metric is defined as the average of the speed percentage over all checks.
    Driving $\text{Driving Efficiency}=\frac{\sum_{i}\text{Speed Percentage}\phantom{}_{i}}{\text{Speed Check Times}}$ (3)
    If the ego vehicle fails to pass the initial 5% checkpoint, this route is not included in the final driving efficiency metric calculation. To account for cases where abnormal speed spikes may occur (e.g., when the vehicle falls off the current map layer), speed percentage values exceeding 1000% are filtered out.
    Comfortness: Comfortness is closely related to human experience and thus requires comparing autonomous driving policy with the behavior of numerous human driving experts to measure it. For this, we follow the popular benchmark nuPlan’s [38] smoothness(also called comfort) protocol, which evaluates ego’s minimum and maximum longitudinal accelerations, the maximum absolute values of lateral acceleration, yaw rate, yaw acceleration, the longitudinal component of jerk, and the maximum magnitude of the jerk vector. These variables are compared to thresholds with default values determined empirically from the examination of nuPlan’s human expert trajectories. Comfortness is measured based on whether these values fall within the upper and lower bounds of the expert values.
    $\begin{split}\text{Frame Variable Smoothness (FVS)}&=\begin{cases}\text{True}&\text{if lower bound}\leq p_{\,i}\leq\text{upper bound},\\ \text{False}&\text{otherwise}\end{cases}\\ &p\in\text{smoothness vars},\,0\leq i\leq\text{total frames}\end{split}$ (4)
    where smoothness variables(vars) include: longitudinal acceleration - expert bound: [-4.05, 2.40], maximum absolute lateral acceleration - expert bound: [-4.89, 4.89], yaw rate - expert bound: [-0.95, 0.95], yaw acceleration - expert bound: [-1.93, 1.93], longitudinal component of jerk - expert bound: [-4.13, 4.13], maximum magnitude of jerk vector - expert bound:[-8.37, 8.37].
    A trajectory is deemed Smooth only if all smoothness variables meet the smoothness criteria.
    Trajectory $\text{Trajectory Smoothness}=\bigwedge_{i=0}^{\text{total frames}}\text{FVS}$
    In nuPlan, smoothness is determined by frame-by-frame evaluation of these variables over the entire trajectory, which makes it susceptible to local driving behaviors. For example, if a vehicle ahead suddenly brakes, the ego vehicle must also brake abruptly to avoid a collision. Even if the ego’s hard brake behavior is appropriate in this case and its driving is smooth at other times, the entire trajectory could still be judged as unsmooth, leading to unreasonable evaluation results. To mitigate this issue, we segment the entire trajectory at a timestep interval $n=20$ for evaluation.
    Segment $\text{Segment Smoothness}=\bigwedge_{i=\text{start frame}}^{\text{end frame}}\text{FVS}$
    The final smoothness metric is defined as the ratio of smooth trajectory segments to the total number of segments.
    Smoothness=Number of Smoothness SegmentsTotal $\text{Smoothness}=\frac{\text{Number of Smoothness Segments}}{\text{Total Segments}}$
    Specifically, if the ego vehicle is blocked (speed remains below 0.1 for more than 60 seconds.), resulting in a failure case, this segment is still be considered as smooth because its speed is safe for human. Note that if the total frames of a trajectory are less than 20, the respective route is excluded from the smoothness assessment.
@@ 4.1 Baselines & Datasets
Table 3: Open-loop and Closed-loop Results of E2E-AD Methods in Bench2Drive under base training set. Avg. L2 is averaged over the predictions in 2 seconds under 2Hz, similar to UniAD. * denotes expert feature distillation.
Method Open-loop Metric Closed-loop Metric
Avg. L2 $\downarrow$ Driving Score $\uparrow$ Success Rate(%) $\uparrow$ Efficiency $\uparrow$ Comfortness $\uparrow$
AD-MLP [23] 3.64 18.05 0.00 48.45 22.63
UniAD-Tiny [4] 0.80 40.73 13.18 123.92 47.04
UniAD-Base [4] 0.73 45.81 16.36 129.21 43.58
VAD [22] 0.91 42.35 15.00 157.94 46.01
TCP* [5] 1.70 40.70 15.00 54.26 47.80
TCP-ctrl* - 30.47 7.27 55.97 51.51
TCP-traj* 1.70 59.90 30.00 76.54 18.08
TCP-traj w/o distillation 1.96 49.30 20.45 78.78 22.96
ThinkTwice* [30] 0.95 62.44 31.23 69.33 16.22
DriveAdapter* [27] 1.01 64.22 33.08 70.22 16.01
Table 4: Multi-Ability Results of E2E-AD Methods under base training set. * denotes expert feature distillation.
@@
AD-MLP [23] 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Tiny [4] 8.89 9.33 20.00 20.00 15.43 14.73
UniAD-Base [4] 14.10 17.78 21.67 10.00 14.21 15.55
VAD [22] 8.11 24.44 18.64 20.00 19.15 18.07
TCP* [5] 16.18 20.00 20.00 10.00 6.99 14.63
TCP-ctrl* 10.29 4.44 10.00 10.00 6.45 8.23
TCP-traj* 8.89 24.29 51.67 40.00 46.28 34.22
TCP-traj w/o distillation 17.14 6.67 40.00 50.00 28.72 28.51
ThinkTwice* [30] 27.38 18.42 35.82 50.00 54.23 37.17
DriveAdapter* [27] 28.82 26.38 48.76 50.00 56.43 42.08
@@ 4.2 Results
Open-loop metric could indicate model convergence but it fails for advanced comparison.. AD-MLP has a high L2 error and performs extremely bad in closed-loop evaluation while VAD has a low L2 error and a decent closed-loop performance. It shows that we could use L2 error to verify the convergence and fitting status of neural networks, i.e., when the L2 error is very high, there should be something wrong within the system. In this case, AD-MLP does not use raw sensors, which is similar to drive blindly and thus infeasible to fit the dataset. Notably, different from findings in nuScenes [21], AD-MLP fails to achieve decent L2 error in Bench2Drive, due to the better behavior diversity as shown in Fig. 5. On the other hand, UniAD-base has a lower L2 error compared to VAD yet with worse closed-loop performance, aligning with findings in [19, 24]. Open-loop evaluation ignore the issues including distribution shift [25] and causal confusion [26, 27] and thus fails to give meaningful comparsion for models with good fitting of dataset, demonstrating the importance of closed-loop evaluation. For efficiency and smoothness, we could observe that AD-MLP has the lowest efficiency due to its quick failure and stuck. UniAD has higher efficiency and smoother trajectories compared to TCP-traj, demonstrating the effectivenes of the UniAD’s post optimization for the planning head.
Expert feature distillation offers important guidance. As pointed out in [49, 50], due to the high-dimensional input space of AD, i.e., multiple images and point clouds, E2E-AD methods tend to overfit. The features from expert, which already possesses strong driving knowledge, could be helpful to mitigate the issue by distillation. As a result, methods (TCP/ThinkTwice/DriveAdapter) with expert feature distillation outperforms those without (VAD/UniAD) by a large margin. From the comparison between TCP-traj with and without distillation, we could observe similar trend. However, in the real world setting, it could be difficult to obtain expert features, which worths further study.
@@
4.3 Case Analysis
We conduct visualizations and upload the results to https://github.com/Thinklab-SJTU/Bench2DriveZoo/blob/uniad/vad/analysis/analysis.md. For all five abilities, we choose some representative scenarios to visualize, where some baselines success and some baselines fail for the ease of comparison and analysis. We give the corresponding failure analysis so that the users and practioners could have a sense about the pros, cons, and future works of existing E2E-AD methods.